OpenStudy (anonymous):

For the system of linear equations below, use graphing to determine in which quadrant the solution lies. x+2y=-2 -2x+y=14

OpenStudy (anonymous):

Quadrant I Quadrant II Quadrant III Quadrant IV

OpenStudy (anonymous):

solving simultaneously y=2 and x therefore equals -6. therefore the solution is in the top left quadrant. (since y is positive but x is negative). solving t grapgically would just involve rearanging them both into the form y=mx +c and drwing them and seeing where they overlap. that would give y=-x/2 -1 and y=2x+14